In this multi-part problem, we will consider this system of simultaneous equations:
3x+5y-6z =2,
5xy-10yz-6xz = -41,
xyz=6.

Let a=3x, b=5y, and c=-6z.
Determine the monic cubic polynomial in terms of a variable t whose roots are t=a, t=b, and t=c. Make sure to enter your answer in terms of t and only t, in expanded form.

Answer by ikleyn(52778)

.
In this multi-part problem, we will consider this system of simultaneous equations:
3x+5y-6z =2,
5xy-10yz-6xz = -41,
xyz=6.

Let a=3x, b=5y, and c=-6z.
Determine the monic cubic polynomial in terms of a variable t whose roots are t=a, t=b, and t=c.
Make sure to enter your answer in terms of t and only t, in expanded form.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

3x  +  5y  - 6z  =   2,       (1)
5xy - 10yz - 6xz = -41,       (2)
xyz = 6                       (3)


Now multiply the equation (2) by 3 (both sides), and multiply the equation (3) by -3*5*6 = -90 (both sides). 
You will get an equivalent system

3x   +  5y  - 6z   =   2,     (1')
15xy - 30yz - 18xz = -41*3    (2')
-90xyz = 6*(-90)              (3')


It is the same as 

3x        +  5y        - 6z         =   2,      (1'')
(3x)*(5y) + (5y)*(-6z) + (3x)*(-6z) = -123      (2'')
(3x)*(5y)*(-6z)                     = -540      (3'')


Using variables a, b and c, you can rewrite the last system equivalently as

a + b + c = 2,                                  (1''')
ab + bc + ac = - 123,                           (2''')
abc = -540.                                     (3''')


Now the monic cubic polynomial in terms of a variable t whose roots are t=a, t=b, and t=c IS

t^3 - 2t^2 + (-123)t - (-540) = 0     ( <<<---+++ as the Vieta's theorem states (!) )

or, which is the same,


t^3 - 2t^2 - 123t + 540 = 0.

Solved.